OpenStudy (elise_a18):

Use the Associative Property of Addition to find the total of 1, 10, and 18 in two different ways. (1 + 10) + 18 = 11 + 18 = 29; (10 + 1) + 18 = 11 + 18 = 29 1 + 10 + 18 = 11 + 18 = 29; 18 + 10 + 1 = 28 + 1 = 29 1 + 10 + 18 = 11 + 18 = 29; 10 + 18 + 1 = 28 + 1 = 29 (1 + 10) + 18 = 11 + 18 = 29; 1 + (10 + 18) = 1 + 28 = 29

OpenStudy (anonymous):

i think its b

OpenStudy (elise_a18):

keep in mind i only give medals to people who EXPLAIN how they got their answer

OpenStudy (anonymous):

ok hold on i need to look up in my book

OpenStudy (chris911):

im going to say C because refer to this http://openstudy.com/study#/updates/503970bce4b09b05dfe3661b

OpenStudy (elise_a18):

omg that's weird... XD

OpenStudy (anonymous):

Test Calculators and Practice Answers archive Word Problems Lessons In Depth Click here to see ALL problems on test Question 790546: Use the Associative Property of Addition to find the total of 1, 10, and 18 in two different ways. A. 1 + 10 + 18 = 11 + 18 = 29; 10 + 18 + 1 = 28 + 1 = 29 B. 1 + 10 + 18 = 11 + 18 = 29; 18 + 10 + 1 = 28 + 1 = 29 C. (1 + 10) + 18 = 11 + 18 = 29; 1 + (10 + 18) = 1 + 28 = 29 D. (1 + 10) + 18 = 11 + 18 = 29; (10 + 1) + 18 = 11 + 18 = 29 (Scroll Down for Answer!) Did you know that Algebra.Com has hundreds of free volunteer tutors who help people with math homework? Anyone can ask a math question, and most questions get answers! OR get immediate PAID help on: Go!!! Answer by solver91311(18808) About Me (Show Source): You can put this solution on YOUR website! Two of the answers move order of the numbers, that is "commutes" them. Think about "commuting to work"; you go from place to place. Two of the answers group the numbers, that is, they cause certain numbers to "associate" with each other in pairs. Find one of the answers that causes the numbers to associate in DIFFERENT ways. John

OpenStudy (chris911):

u got that from this web^ http://www.algebra.com/algebra/homework/playground/test.faq.question.790546.html

OpenStudy (anonymous):

so

OpenStudy (anonymous):

i didnt fine it in the book

OpenStudy (chris911):

but they didn't give the answer these website goive the answer lol http://openstudy.com/study#/updates/503970bce4b09b05dfe3661b and http://www.justanswer.com/math-homework/7e4j4-need-help-homework-please.html

OpenStudy (anonymous):

oh it said A

OpenStudy (chris911):

I have 2 places where it says C so Elise u make the decision

OpenStudy (anonymous):

kk